# Trees of Graphs as Boundaries of Hyperbolic Groups

### v1

@@ 6. Regular trees of graphs
Let ${\mathcal{N}}_{\Gamma}$ be a standard family of normal neighbourhoods of $\Gamma$, as described in Example 2.2, and let $\Gamma^{\circ}:=|\Gamma|\setminus\bigcup\{\operatorname{int}B:B\in{\mathcal{N}}_{\Gamma}\}$ be the space with peripherals, called the punctured graph $\Gamma$, as described in the same example. Denote by $\Omega^{E}_{\Gamma}=\{{\rm bd}(B):B\in{\mathcal{N}}_{\Gamma},B\cap V_{\Gamma}=\emptyset\}$ the set of those peripherals in $\Gamma^{\circ}$ which correspond to those $B\in{\mathcal{N}}_{\Gamma}$ which are normal neighbourhoods of interior points of the edges of $\Gamma$. Similarly, put $\Omega_{\Gamma}^{V}=\{{\rm bd}(B):B\in{\mathcal{N}}_{\Gamma},B\cap V_{\Gamma}\neq\emptyset\}$. We will call the elements of the above sets E-peripherals and V-peripherals, respectively. Observe that each E-peripheral is a set of cardinality 2, i.e. a doubleton. Each such set can be equipped with one of the two possible orders, so that it becomes an ordered pair of points, and we will call any such order an orientation associated to this E-peripheral. We denote by $\widetilde{\Omega}^{E}_{\Gamma}$ the set of all oriented E-peripherals of $\Gamma^{\circ}$, i.e. the set $\{(p,q):\{p,q\}\in\Omega^{E}_{\Gamma}\}$. Observe that each of the oriented E-peripherals may be either compatible or incompatible with the fixed orientation of the edge $e$ of $\Gamma$ in the interior of which it is contained. For any oriented peripheral $\delta\in\widetilde{\Omega}^{E}_{\Gamma}$, we denote by $|\delta|$ the corresponding underlying (unoriented) peripheral, and by $\bar{\delta}$ the same peripheral $|\delta|$ oriented oppositely than $\delta$.
We will now describe some auxiliary labelling $\lambda$ of the family $\widetilde{\Omega}_{\Gamma}^{E}$. The labels of this labelling will be in the set $O_{\Gamma}$ of all oriented edges of $\Gamma$. We use the notation $|\varepsilon|$ and $\bar{\varepsilon}$, respectively, for the underlying unoriented edge of an oriented edge $\varepsilon$, and for the oppositely oriented edge. For any oriented edge $\varepsilon$ of $\Gamma$, consider the set
$O_{\varepsilon}=\{\varepsilon^{\prime}:(\varepsilon,\varepsilon^{\prime})\in{\mathcal{A}}\}.$
By condition (3) of Definition 6.1, this set is is nonempty. Take as $\lambda$ any labelling of $\widetilde{\Omega}_{\Gamma}^{E}$ such that:
@@
    for any $\varepsilon\in O_{\Gamma}$, and for any oriented peripheral $\delta$ contained in the interior of $|\varepsilon|$ whose orientation is compatible with that of $\varepsilon$, the label of $\delta$ is in the subset $O_{\varepsilon}$, i.e. $\lambda(\delta)\in O_{\varepsilon}$;
@@
    $\{|\delta|\in\Omega^{E}_{\Gamma}:\text{$|\delta|\subset\operatorname{int}|\varepsilon|,\delta$ is compatible with $\varepsilon$ and $\lambda(\delta)=\varepsilon^{\prime}$}\}$
@@
A labelling $\lambda$ as above clearly exists, and it is unique in the sense of the following claim, a straightforward proof of which we omit.
@@
For any two labellings $\lambda,\lambda^{\prime}$ as above, satisfying conditions (L1)(L3), there is a homeomorphism $h:|\Gamma|\to|\Gamma|$ which fixes all vertices of $\Gamma$, which preserves the family ${\mathcal{N}}_{\Gamma}$, and whose restriction $h^{\circ}=h|_{\Gamma^{\circ}}$ is a homeomorphism of $\Gamma^{\circ}$ such that $\lambda^{\prime}(h^{\circ}(\delta)=\lambda(\delta)$ for any $\delta\in\widetilde{\Omega}_{\Gamma}^{E}$ (where we use here a convention that $h^{\circ}$ maps oriented E-peripherals to oriented E-peripherals). Moreover, if $\delta,\delta^{\prime}\in\widetilde{\Omega}^{E}_{\Gamma}$ are contained in the interior of the same edge of $\Gamma$, if they are consistently oriented, and if $\lambda(\delta)=\lambda^{\prime}(\delta^{\prime})$, then $h$ above can be chosen so that $h^{\circ}(\delta)=\delta^{\prime}$.
We are now ready to describe the tree system $\Theta[{\mathcal{R}}]=(T,\{\Gamma_{t}^{\circ}\},\{\Sigma_{u}\},\{\varphi_{e}\})$. As the underlying tree $T$ take the unique bipartite tree, with the vertex set partitioned into black and white vertices, such that the order of any white vertex is equal to 2, and order of any black vertex is countable infinite. Now, we describe resursively the punctured graphs $\Gamma_{t}^{\circ}$ associated to the black vertices of $T$. Let $t_{0}$ be any black vertex of $T$ (which we will view as the initial vertex). Put $\Gamma_{t_{0}}$ to be any connected component $\Gamma_{i}$ of $\Gamma$. Fix also some auxiliary bijection $\beta_{t_{0}}$ between the set $N_{t_{0}}$ of all (white) vertices of $T$ adjacent to $t_{0}$ and the set of peripherals in the punctured graph $\Gamma_{t_{0}}^{\circ}$. (We will be choosing appropriate bijections $\beta_{t}:t\in V^{b}_{T}$ simultaneously with the description of the constituent spaces $\Gamma_{t}^{\circ}$; these bijections will play more essential role at later stages of the construction of the tree systrem $\Theta[{\mathcal{R}}]$.) Now, for each $k\geq 0$, denote by $T_{k}$ the b-subtree of $T$ spanned by all vertices of $T$ lying at polygonal distance $\leq 2k$ from $t_{0}$. Suppose that for all $t\in V^{b}_{T_{k}}$ we have already chosen the punctured graphs $\Gamma_{t}^{\circ}$ and the corresponding bijections $\beta_{t}$ from the sets $N_{t}$ (of all neighbours of $t$ in the tree $T$) to the sets of peripherals in $\Gamma_{t}^{\circ}$, so that the following conditions hold:
@@
      if $\beta_{t}(u)$ is a V-peripheral, i.e. $\beta_{t}(u)={\rm bd}(B)$ for $B$ containing a vertex of $\Gamma_{i(t)}$, say $v$, then $\beta_{t^{\prime}}(u)$ is also a V-peripheral; moreover, if $v^{\prime}$ is the corresponding vertex of $\Gamma_{i(t^{\prime})}$ for $\beta_{t^{\prime}}(u)$, then, viewing $v,v^{\prime}$ naturally as vertices of $\Gamma$, we have $a(v)=v^{\prime}$.
@@
As the main recursive step of the description of constituent spaces, we describe now the spaces $\Gamma_{t}^{\circ}$, and the associated bijections $\beta_{t}$, for all vertices $t\in V^{b}_{T_{k+1}}\setminus V^{b}_{T_{k}}$, so that the above conditions (R1) and (R2) will be satisfied with $k$ replaced by $k+1$. Consider any vertex $t\in V^{b}_{T_{k+1}}\setminus V^{b}_{T_{k}}$, and denote by $s$ the unique vertex in $V^{b}_{T_{k}}$ at combinatorial distance 2 from $t$, and by $u$ the unique common neighbour of both $t$ and $s$ (which is a white vertex). Suppose first that the subspace $\beta_{s}(u)\subset\Gamma_{s}^{\circ}$ is an E-peripheral, i.e. it is contained in the interior of a single edge, say $|\varepsilon|$, of $\Gamma_{s}$, and denote by $\delta$ this same peripheral oriented consistently with $\varepsilon$. Recall that, under our identification conventions, the oriented periheral $\delta$ has a label, and denote this label by $\varepsilon^{\prime}$. Put $\Gamma_{t}$ to be this component of $\Gamma$ which contains $|\varepsilon^{\prime}|$, and denote this component by $\Gamma_{i(t)}$. Take as $\beta_{t}$ any bijection from $N_{t}$ to the set of peripherals of $\Gamma_{t}^{\circ}$ for which $\beta_{t}(u)$ is contained in the interior of $|\varepsilon^{\prime}|$ and, after orienting it consistetly with $\varepsilon^{\prime}$, it is labeled with $\varepsilon$. Suppose now that $\beta_{s}(u)$ is a V-peripheral, and denote by $v$ this vertex of $\Gamma_{s}$ for which $\beta_{s}(u)=\partial B$, where $B$ is a normal neighbourhood of $v$. Identifying $\Gamma_{s}$ with the component $\Gamma_{i(s)}$ of $\Gamma$, put $v^{\prime}:=a(v)$. Denote by $\Gamma_{i(t)}$ this component of $\Gamma$ which contains $v^{\prime}$, and put $\Gamma_{t}=\Gamma_{i(t)}$. Finally, take as $\beta_{t}$ any bijection from $N_{t}$ to the set of peripherals of $\Gamma_{t}^{\circ}$ for which $\beta_{t}(u)=\partial B^{\prime}$, where $B^{\prime}$ is a normal neighbourhood of $v^{\prime}$. We skip a straightforward verification that this provides data as required.
@@
We now turn to describing the sets $\Sigma_{u}$ and the maps $\varphi_{e}$ of $\Theta[{\mathcal{R}}]$. Let $u$ be any white vertex of $T$, and denote by $t,s$ the two black vertices adjacent to $u$. Suppose first that $\beta_{t}(u)$ is an E-peripheral of $\Gamma_{t}^{\circ}$, i.e. it is contained in the interior of some edge, say $|\varepsilon|$, of $\Gamma_{t}$. Observe that, by the construction, the peripheral $\beta_{s}(u)$ is then also contained in the interior of an edge of the corresponding graph $\Gamma_{s}$, and denote this edge by $|\varepsilon^{\prime}|$. Choose also the orientations of the peripherals $\beta_{t}(u)$ and $\beta_{s}(u)$, getting the oriented peripherals $\delta$ and $\delta^{\prime}$, respectively, and assume without loss of generality (changing $\varepsilon^{\prime}$ into $\overline{\varepsilon^{\prime}}$, if necessary) that $\lambda(\delta)=\varepsilon^{\prime}$ and $\lambda(\delta^{\prime})=\varepsilon$. Put $\Sigma_{u}$ to be a set consisting of two points, with discrete topology, and consider an auxiliary order on this set. Put $\varphi_{[t,u]}:\Sigma_{u}\to|\delta|$ and $\varphi_{[s,u]}:\Sigma_{u}\to|\delta^{\prime}|$ to be the bijections which preserve the distinguished orders.
@@
Finally, suppose that $\beta_{t}(u)$ is a V-peripheral, and denote by $v$ this vertex of $\Gamma_{t}$ for which $\beta_{t}(u)=\partial B$, where $B$ is a normal neighbourhood of $v$. Similarly, denote by $v^{\prime}$ this vertex of $\Gamma_{s}$ for which $\beta_{s}(u)=\partial B^{\prime}$, where $B^{\prime}$ is a normal neighbourhood of $v^{\prime}$. Note that, by R2(b), we have $a(v)=v^{\prime}$, and that, identifying $\Gamma_{t}$ and $\Gamma_{s}$ as the corresponding components of $\Gamma$, we have a bijection $\bar{\alpha}_{v}:\beta_{t}(u)\to\beta_{s}(u)$ naturally induced by the bijection $\alpha_{v}$. Take as $\Sigma_{u}$ any set of the same cardinality as $\beta_{t}(u)$ (and as $\beta_{s}(u)$), with discrete topology. Take as the maps $\varphi_{[u,t]}$ and $\varphi_{[u,s]}$ any bijections for which the composition $\varphi_{[u,s]}\circ\varphi_{[u,t]}^{-1}$ coincides with $\bar{\alpha}_{v}$.
@@
For any finite connected graph $\Gamma$, let ${\mathcal{R}}^{r}_{\Gamma}=(\Gamma,\mathbf{a},\mathcal{A})$ be a graphical connecting system in which $a={\rm id}_{V_{\Gamma}}$, $\alpha_{v}={\rm id}_{\operatorname{Lk}(v)}$ for each $v\in V_{\Gamma}$, and ${\mathcal{A}}=\{(\varepsilon,\varepsilon):\varepsilon\in O_{\Gamma}\}$, where $O_{\Gamma}$ denotes the set of all oriented edges of $\Gamma$. We call ${\mathcal{R}}^{r}_{\Gamma}$ the reflection connecting system for $\Gamma$. The associated tree systems $\Theta[{\mathcal{R}}^{r}_{\Gamma}]$ lead to a class of regular trees of graphs which are called reflection trees of graphs $\Gamma$ and which have been earlier described and studied in [24].
@@
A regular tree of graphs is any space of form ${\mathcal{X}}({\mathcal{R}})$, where $\mathcal{R}$ is a graphical connecting system. We also call the space ${\mathcal{X}}({\mathcal{R}})$ the regular tree of graphs for $\mathcal{R}$.
@@
Note that, if $\mathcal{R}$ is 2-connected, then obviously the corresponding system $\Theta[{\mathcal{R}}]$ is a tree system of punctured 2-connected graphs, and consequently, the regular tree of graphs ${\mathcal{X}}(\mathcal{R})$ is then connected, locally connected, and has topological dimension 1 (see Lemma 4.14 and Lemma 4.15 above). We call the spaces ${\mathcal{X}}(\mathcal{R})$ as above the regular trees of 2-connected graphs.

### v2

@@ 6. Regular trees of graphs
Recall that we denote by $\widetilde{\Omega}_{\Gamma}^{E}$ the set of all oriented E-peripherals in the punctured graph $\Gamma^{\circ}$. We will now describe some auxiliary labelling $\lambda$ of the family $\widetilde{\Omega}_{\Gamma}^{E}$. The labels of this labelling will be in the set $O_{\Gamma}$ of all oriented edges of $\Gamma$. Observe that, by condition (3) of Definition 6.1, for any oriented edge $\varepsilon$ of $\Gamma$ the corresponding set $O_{\varepsilon}$ (as defined in the statement of condition (4) of Definition 6.2) is nonempty. Take as $\lambda$ any labelling of $\widetilde{\Omega}_{\Gamma}^{E}$ such that:
@@
    for any $\varepsilon\in O_{\Gamma}$, and for any oriented peripheral $\delta$ contained in the interior of $|\varepsilon|$ whose orientation is compatible with that of $\varepsilon$, the label of $\delta$ is in the subset $O_{\varepsilon}$, i.e., $\lambda(\delta)\in O_{\varepsilon}$;
@@
    $\{|\delta|\in\Omega^{E}_{\Gamma}:\text{$|\delta|\subset\operatorname{int}|\varepsilon|,\delta$ is oriented compatibly with $\varepsilon$, and $\lambda(\delta)=\varepsilon^{\prime}$}\}$
@@
It is not hard to realize (and we omit the details) that a labelling $\lambda$ as above exists, and is unique in the sense of the first part of the following claim. We omit the proof, which is a standard back-and-forth argument.
@@
For any two labellings $\lambda,\lambda^{\prime}$ as above, satisfying conditions (L1)(L3), there is a homeomorphism $h:|\Gamma|\to|\Gamma|$ which fixes all vertices of $\Gamma$, which preserves all edges of $\Gamma$, which preserves the family ${\mathcal{N}}_{\Gamma}$, and whose restriction $h^{\circ}=h|_{\Gamma^{\circ}}$ is a homeomorphism of $\Gamma^{\circ}$ such that $\lambda^{\prime}(h^{\circ}(\delta))=\lambda(\delta)$ for any $\delta\in\widetilde{\Omega}_{\Gamma}^{E}$ (where $h^{\circ}(\delta)$ is viewed as an oriented E-peripheral with orientation induced from $\delta$ via $h^{\circ}$). Moreover, if $\delta,\delta^{\prime}\in\widetilde{\Omega}^{E}_{\Gamma}$ are contained in the interior of the same edge of $\Gamma$, if they are consistently oriented, and if $\lambda(\delta)=\lambda^{\prime}(\delta^{\prime})$, then $h$ above can be chosen so that $h^{\circ}(\delta)=\delta^{\prime}$.
We are now ready to describe the tree system $\Theta[{\mathcal{R}}]=(T,\{\Gamma_{t}^{\circ}\},\{\Sigma_{u}\},\{\varphi_{e}\})$. As the underlying tree $T$ take the unique bipartite tree, with the vertex set partitioned into black and white vertices, such that the degree of any white vertex is equal to 2, and the degree of any black vertex is countable infinite. Now, we describe resursively the punctured graphs $\Gamma_{t}^{\circ}$ associated to the black vertices of $T$. Let $t_{0}$ be any black vertex of $T$ (which we will view as the initial vertex). Put $\Gamma_{t_{0}}$ to be any connected component $\Gamma_{i}$ of $\Gamma$. Fix also some auxiliary bijection $\beta_{t_{0}}$ between the set $N_{t_{0}}$ of all (white) vertices of $T$ adjacent to $t_{0}$ and the set of peripherals in the punctured graph $\Gamma_{t_{0}}^{\circ}$. (We will be choosing appropriate bijections $\beta_{t}:t\in V^{b}_{T}$ simultaneously with the description of the constituent spaces $\Gamma_{t}^{\circ}$; these bijections will play an essential role at later stages of the construction of the tree system $\Theta[{\mathcal{R}}]$.) Now, for each $k\geq 0$, denote by $T_{k}$ the b-subtree of $T$ spanned by all vertices of $T$ lying at distance $\leq 2k$ from $t_{0}$. Suppose that for all $t\in V^{b}_{T_{k}}$ we have already chosen the punctured graphs $\Gamma_{t}^{\circ}$ and the corresponding bijections $\beta_{t}$ from the sets $N_{t}$ (of all neighbours of $t$ in the tree $T$) to the sets of peripherals in $\Gamma_{t}^{\circ}$, so that the following conditions hold:
@@
      if $\beta_{t}(u)$ is a V-peripheral, i.e., $\beta_{t}(u)={\rm bd}(B)$ for $B$ a normal neighborhood of a vertex $v$ of $\Gamma_{i(t)}$ then $\beta_{t^{\prime}}(u)$ is also a V-peripheral; moreover, if $v^{\prime}$ is the corresponding vertex of $\Gamma_{i(t^{\prime})}$ for $\beta_{t^{\prime}}(u)$, then, viewing $v,v^{\prime}$ as vertices of $\Gamma$, we have $a(v)=v^{\prime}$.
@@
As the main recursive step of the description of constituent spaces, we describe now the spaces $\Gamma_{t}^{\circ}$, and the associated bijections $\beta_{t}$, for all vertices $t\in V^{b}_{T_{k+1}}\setminus V^{b}_{T_{k}}$, so that the above conditions (R1) and (R2) will be satisfied with $k$ replaced by $k+1$. Consider any vertex $t\in V^{b}_{T_{k+1}}\setminus V^{b}_{T_{k}}$, and denote by $s$ the unique vertex in $V^{b}_{T_{k}}$ at combinatorial distance 2 from $t$, and by $u$ the unique common neighbour of both $t$ and $s$ (which is a white vertex). Suppose first that the subspace $\beta_{s}(u)\subset\Gamma_{s}^{\circ}$ is an E-peripheral, i.e., it is contained in the interior of a single edge, say $|\varepsilon|$, of $\Gamma_{s}$, and denote by $\delta$ this same peripheral oriented consistently with $\varepsilon$. Recall that, under our identification conventions, the oriented peripheral $\delta$ has a label, and denote this label by $\varepsilon^{\prime}$. Put $\Gamma_{t}$ to be a copy of this component of $\Gamma$ which contains $|\varepsilon^{\prime}|$, and denote this component by $\Gamma_{i(t)}$. Take as $\beta_{t}$ any bijection from $N_{t}$ to the set of peripherals of $\Gamma_{t}^{\circ}$ for which $\beta_{t}(u)$ is contained in the interior of $|\varepsilon^{\prime}|$ and, after orienting it consistently with $\varepsilon^{\prime}$, it is labeled with $\varepsilon$. Suppose now that $\beta_{s}(u)$ is a V-peripheral, and denote by $v$ this vertex of $\Gamma_{s}$ for which $\beta_{s}(u)=\partial B$, where $B$ is a normal neighbourhood of $v$. Identifying $\Gamma_{s}$ with the component $\Gamma_{i(s)}$ of $\Gamma$, put $v^{\prime}:=a(v)$. Denote by $\Gamma_{i(t)}$ this component of $\Gamma$ which contains $v^{\prime}$, and put $\Gamma_{t}=\Gamma_{i(t)}$. Finally, take as $\beta_{t}$ any bijection from $N_{t}$ to the set of peripherals of $\Gamma_{t}^{\circ}$ for which $\beta_{t}(u)=\partial B^{\prime}$, where $B^{\prime}$ is a normal neighbourhood of $v^{\prime}$. We skip a straightforward verification that this provides data as required.
@@
We now turn to describing the sets $\Sigma_{u}$ and the maps $\varphi_{e}$ of $\Theta[{\mathcal{R}}]$. Let $u$ be any white vertex of $T$, and denote by $t,s$ the two black vertices adjacent to $u$. Suppose first that $\beta_{t}(u)$ is an E-peripheral of $\Gamma_{t}^{\circ}$, i.e., it is contained in the interior of some edge, say $|\varepsilon|$, of $\Gamma_{t}$. Observe that, by the construction, the peripheral $\beta_{s}(u)$ is then also contained in the interior of an edge of the corresponding graph $\Gamma_{s}$, and denote this edge by $|\varepsilon^{\prime}|$. Then $\varepsilon$ and $\varepsilon^{\prime}$ induce orientations on $\beta_{t}(u)$ and $\beta_{s}(u)$ giving us oriented E-peripherals $\delta$ and $\delta^{\prime}$, respectively. Without loss of generality (changing $\varepsilon^{\prime}$ into $\overline{\varepsilon^{\prime}}$ and $\delta^{\prime}$ into $\overline{\delta^{\prime}}$, if necessary) we have $\lambda(\delta)=\varepsilon^{\prime}$ and $\lambda(\delta^{\prime})=\varepsilon$. Put $\Sigma_{u}$ to be a set consisting of two points, with discrete topology, and consider an auxiliary order on this set. Put $\varphi_{[t,u]}:\Sigma_{u}\to|\delta|$ and $\varphi_{[s,u]}:\Sigma_{u}\to|\delta^{\prime}|$ to be the bijections which preserve the distinguished orders.
@@
Finally, suppose that $\beta_{t}(u)$ is a V-peripheral, and denote by $v$ this vertex of $\Gamma_{t}$ for which $\beta_{t}(u)=\partial B$, where $B$ is a normal neighbourhood of $v$. Similarly, denote by $v^{\prime}$ this vertex of $\Gamma_{s}$ for which $\beta_{s}(u)=\partial B^{\prime}$, where $B^{\prime}$ is a normal neighbourhood of $v^{\prime}$. Note that, by (R2)(b), we have $a(v)=v^{\prime}$, and that, identifying $\Gamma_{t}$ and $\Gamma_{s}$ with the corresponding components of $\Gamma$, we have a bijection $\hat{\alpha}_{v}:\beta_{t}(u)\to\beta_{s}(u)$ naturally induced by the bijection $\alpha_{v}$. Take as $\Sigma_{u}$ any set of the same cardinality as $\beta_{t}(u)$ (and as $\beta_{s}(u)$), with discrete topology. Take as the maps $\varphi_{[u,t]}$ and $\varphi_{[u,s]}$ any bijections for which the composition $\varphi_{[u,s]}\circ\varphi_{[u,t]}^{-1}$ coincides with $\hat{\alpha}_{v}$.
@@
For any finite connected graph $\Gamma$, let ${\mathcal{R}}^{r}_{\Gamma}=(\Gamma,\mathbf{a},\mathcal{A})$ be a graphical connecting system in which $a={\rm id}_{V_{\Gamma}}$, $\alpha_{v}={\rm id}_{\operatorname{Lk}(v)}$ for each $v\in V_{\Gamma}$, and ${\mathcal{A}}=\{(\varepsilon,\varepsilon):\varepsilon\in O_{\Gamma}\}$, where $O_{\Gamma}$ denotes the set of all oriented edges of $\Gamma$. We call ${\mathcal{R}}^{r}_{\Gamma}$ the reflection connecting system for $\Gamma$. The associated tree systems $\Theta[{\mathcal{R}}^{r}_{\Gamma}]$ lead to a class of regular trees of graphs (see Definition 6.6 below) which are called reflection trees of graphs and which have been described and studied earlier in [Ś21].
@@
A regular tree of graphs is any space of the form ${\mathcal{X}}({\mathcal{R}})=\lim\Theta[{\mathcal{R}}]$, where $\mathcal{R}$ is a graphical connecting system, and where $\Theta[{\mathcal{R}}]$ is the (unique up to an isomorphism) tree system of punctured graphs compatible with $\mathcal{R}$. Moreover, we call any space ${\mathcal{X}}({\mathcal{R}})$ as above the regular tree of graphs for $\mathcal{R}$.
@@
Note that, if $\mathcal{R}$ is 2-connected, then obviously the corresponding system $\Theta[{\mathcal{R}}]$ is a tree system of punctured 2-connected graphs, and consequently, the regular tree of graphs ${\mathcal{X}}(\mathcal{R})$ is then connected, locally connected, cutpoint-free and has topological dimension 1 (see Lemma 4.14 and Lemma 4.15 above). We call the spaces ${\mathcal{X}}(\mathcal{R})$ as above the regular trees of 2-connected graphs.